OpenStudy (anonymous):

Can anyone help? f(x)=x^2-x-2/2x^2-x-21 I need to find hortizontal, vertical and oblique. I couldn't find any vertical/oblique, but for Hortizontal I have x-3 and x=7/2

OpenStudy (anonymous):

degree of numerator is equal to the degree of the denominator so there is no oblique asymptote

OpenStudy (anonymous):

but since they are equal the horizontal asymptote is the ratio of the leading coefficients, in your case it is \(y=\frac{1}{2}\)

OpenStudy (anonymous):

how do i set the problem up?

OpenStudy (anonymous):

i am not sure what you mean by "set the problem up" it is a think problem degree of numerator is 2 degree of denominator is 2 since they are equal, the horizontal asymptote is the ratio of the leading coefficients the leading coefficient of the numerator is 1 the leading coefficient of the denominator is 2 the ratio is therefore \(\frac{1}{2}\) and your horizontal asymptote is \(y=\frac{1}{2}\) you pretty much do it in your head

OpenStudy (anonymous):

for vertical set \(2x^2-x-21=0\) and solve for \(x\)

OpenStudy (anonymous):

which looks like what you did for "horizontal" but it should be vertical you get \[(x+3)(2x-7)=0\] \[x=-3\] or \[x=\frac{7}{2}\]for the vertical asymptotes

OpenStudy (anonymous):

is my hortizontal correct? I took (x-2)(x+1)/(x+3)(2x-7)

OpenStudy (anonymous):

so my hortizontal is x=-3 and vertical is y=1/2